OpenStudy (peace420):

Please help!? What ratio that you would use to convert 7,920 miles to feet? Even if you do not recall the numbers used, be sure to explain in steps how to do the conversion.

OpenStudy (jonathan34):

Decimals?

OpenStudy (sshayer):

1 mile=1760 yards 1 yard=3 feet

OpenStudy (jonathan34):

sshayer would you know a decimal?

OpenStudy (sshayer):

i don't understand why are you mentioning decimals?

OpenStudy (jonathan34):

because with decimals we can find the feet

OpenStudy (sshayer):

first change in yards by above conversion formula. Then in feet

OpenStudy (sshayer):

i am not aware of it.

OpenStudy (sshayer):

1 mile=1760 yards=1760*3=5280 feet 7920 miles=7920*5280=?

OpenStudy (peace420):

thanks @sshayer

OpenStudy (sshayer):

yw

OpenStudy (triciaal):

the question wants you to explain how to change from one unit of measure to the other using the conversion factor. notice "even if you do not recall the number"